OpenStudy (anonymous):

A) Write an Equation B) Solve it The perimeter of the rectangle is 78 feet. Find it's dimensions if the length is 2x-4 and the width is 2x+3 (I'LL FAN AND MEDAL)

OpenStudy (imstuck):

The formula for the perimeter of a rectangle is P = 2L + 2W, right? They give you the length as 2x-4 and the width as 2x+3. They also give you the perimeter as 78. So fitting all that info into our equation, we get 78 = 2(2x-4) + 2(2x+3) 78 = 4x - 8 + 4x + 6 78 = 8x - 2 80 = 8x 10 = x

OpenStudy (imstuck):

Now we use the fact that x = 10 and find the length and the width with it. Like this: L = 2x - 4---> L = 2(10) - 4---> L = 20-4---> L = 16 W = 2x + 3---> W = 2(10) + 3---> W = 20 + 3---> W = 23